Katie Perry earns $5,750 per month working for his employer. The FICA tax rate for Social Security is 6.2% of the first $118,500 earned each calendar year and the FICA tax rate for Medicare is 1.45% of all earnings. The current FUTA tax rate is 0.6%, and the SUTA tax rate is 5.4%. Both unemployment taxes are applied to the first $7,000 of an employee's pay. The employee has $192 in federal income taxes withheld. Katie has voluntary deductions for health insurance of $160 and contributes $80 to a retirement plan each month. What is the amount of net pay for Katie for the month of January? (Round your intermediate calculations to two decimal places.)
Answer #1

FICA tax – social security tax = 5750*6.2% = $356.50

FICA tax – Medicare tax = 5750*1.45% = $83.38

Net Pay = gross pay – social security tax – medicare tax – federal income tax withheld – health insurance deduction – contribution to retirement plan = 5750-356.50-83.38-192-160-80 = $4878.12